4. Evaluate (F•dr for F(x, y, z)=(3xz)i +(2z - y)j+(22)k and с C:r(t)= (2t+1)i+(3).j+(5t)k, Ostsi
Answer #1

? uc f.dr = ? F(X,Y,Z) = (372) + (22-y)+ zak Ble) = (2+++33 +54 k 0 <t<! using the natation ( mdx + + P dz Aldy we have R=(X,Now 1 Su Mda + Ndy + p dz с * = S (301? + 154) 2017 + (10 4-3) (0 ) + 625491544) 0 = s 160ť + 30 t) dt + 125382 dt 1 1 5 608’